OpenStudy (anonymous):

Marcus wants to buy 5 baseballs. He has $35. What is the most each baseball can cost? (I need the inequality for it, please)

OpenStudy (anonymous):

Can the baseballs cost $1? Yes, 5 balls @ $1 each makes $5. But that is not the most they can cost. Can the balls cost $8 each? No, 5 balls @ $8 each would cost $40, more than Marcus has.

OpenStudy (anonymous):

Thanks for the help, but can I please get the inequality form of it?

OpenStudy (anonymous):

so do you know the number, and you're just unsure of the inequality? What is the most they can cost, in plain English? I can help with the inequality, but you need to be able to see the max cost.

OpenStudy (anonymous):

The maximum cost is $7, I'm just unsure of the inequality, like you said.

OpenStudy (anonymous):

ok, cool... you are practically at the answer... the way you would set this up is: cost per ball X number of balls <= amount of money Marcus has cost per ball X 5 <= 35 Then divide both sides by 5: cost per ball <= 7

OpenStudy (anonymous):

\[5 \times cost \le 35\] \[cost \le 7\]

OpenStudy (anonymous):

Oh, I get it now. Thanks a lot! :)

OpenStudy (anonymous):

Glad to help :)